You just purchased 200 shares of a company stock at $50 per share.The dividend is projected to increase at a constant rate of 4.50% per yr. The required rate of
return on the stock is 8.00%. What is the a
stock’s expected price 2 yrs from today? (please use up to four decimal numbers for dividend parts in calculation.)
Answer #1
Dear Student
Thank you for using Chegg
Please find below the answer
Statementshowing Computations
Paticulars Amount
P0 = D1/(ke-g)
Where P0 is the price 50
Ke is the required return 8.00%
g is the growth rate 4.50%
D1 is dividend at end of year ?
50 = D1/(8%-4.5%)
50 = D1/(3.5%)
D1 = $1.75
D3 = 1.75*1.045*1.045                        1.9110
stock’s expected price 2 yrs from today = 1.911 /(8%-4.5%)
stock’s expected price 2 yrs from today = 1.911 /(3.5%)                          54.60